# Inexact Column Generation for Bayesian Network Structure Learning via Difference-of-Submodular Optimization

### v1

@@ 4 Row and Column Generation Scheme
4:     Initialize $\texttt{rc}_{i}\leftarrow-\infty$ for $i=1,\ldots,n$triangleright initialize reduced costs
@@
10:                  $(\texttt{pa},\texttt{rc}_{i})\leftarrow\text{DCA-Pricing}(D,i,x^{0},\lambda^{*},\Lambda,\epsilon)$triangleright DCA pricing
11:                  $\hat{\mathcal{P}}_{i}\leftarrow\hat{\mathcal{P}}_{i}\cup\{\texttt{pa}\}$triangleright column generation
@@
19:              $\hat{\mathcal{C}}\leftarrow\hat{\mathcal{C}}\cup\{C^{*}\}$
@@
22:     $G\leftarrow$ optimal DAG chosen from $\hat{\mathcal{P}}_{1:n}$; simultaneously update $\hat{\mathcal{C}}$triangleright row generation for RMIP
@@
Algorithm 2 demonstrates how we incorporate our DCA pricing method into a simultaneous row and column generation framework. The algorithm begins by initializing the candidate parent sets $\hat{\mathcal{P}}_{1:n}$ and the cluster set $\hat{\mathcal{C}}$ with basic candidates (Line 2). It then iterates among three main phases: (1) a CG phase (Lines 3-14) where the RMLP is solved to obtain optimal dual solutions $\lambda^{*}$, followed by the solution of the pricing problem for each node via DCA — The pricing solutions that have negative reduced costs are added to $\hat{\mathcal{P}}_{i}$. The CG phase repeats until all reduced costs become non-negative. In Line 8, the pricing problem on node $i$ is solved only when the current reduced cost $rc_{i}$ is negative, to avoid spending too much time on pricing; (2) a row generation phase (Lines 15-21) where the RMLP optimal solution is used to identify the most violated cluster constraint (see Section B for details) which is added to cluster set $\hat{\mathcal{C}}$; (3) the integer solution phase (Line 22), where the algorithm finds the optimal DAG $G$ with current candidate $\hat{\mathcal{P}}_{1:n}$ — This is achieved by solving RMIP using branch and cut with callbacks (a feature that allows users to interrupt the solution process in modern IP solvers such as Gurobi), where we check if any cluster constraint is violated when we meet a integer incumbent and add violated cluster constraints as lazy constraints. The three phases iterate until convergence to a valid DAG solution, progressively refining both the columns in $\hat{\mathcal{P}}_{1:n}$ and cluster constraints in $\hat{\mathcal{C}}$.
Slightly different from its description, when we implement the CG phase in Algorithm 2, in addition to the final solution generated by DCA, we add to $\hat{\mathcal{P}}_{i}$ all candidate parent sets associated with intermediate $x^{t}$ solutions that have negative reduced costs. Empirically, we find that it helps accelerate convergence and improves solution quality.
@@ 4.1 Choice of the Initial Point
A step that remains ambiguous in Algorithm 2 is the choice of the initial point $x^{0}$ for DCA (Line 2), which we find having a significant impact on the overall solution time and solution quality. Below we first describe two initialization strategies.
The first strategy is a warm-start initialization strategy. Note that by the complementarity property of optimal solutions in LP bertsimas1997introduction , the parent sets selected (i.e., parent sets $J$ associated with the primal variables $x_{i\leftarrow J}$ taking positive values) in the RMLP solution have reduced costs equal to zero. Therefore, we consider using the parent sets selected in the RMLP solution to construct the initial solution $x^{0}$ for DCA. Due to the monotonicity property (4) of DCA, it often quickly finds columns with negative reduced costs, leading to a rapid convergence of DCA. However, a drawback of the warm-start initialization strategy is that it may restrict exploration of parent sets as the evaluation remains localized around solutions already contained in the candidate parent set choice set. Preliminary experiments show that a pure warm-start initialization strategy can quickly find a DAG with a good score while the solution often deviates a lot from the ground truth due to its greedy nature.
The second strategy employs a random initialization that samples a point $x^{0}$ uniformly from $[0,1]^{V\setminus\{i\}}$. This random initialization strategy facilitates global exploration on the parent set space and may discover patterns distant from the current ones in $\hat{\mathcal{P}}_{i}$. However, we find that exclusive reliance on random initialization can significantly slow down DCA convergence, as computational resources may be expended on evaluating solutions far from the optimum.
A hybrid initialization strategy that combines efficiency and diversity of the previous two strategies is often more effective in practice. In our implementation, we initially employ random initialization to broadly explore the solution space. Once the candidate set $\hat{\mathcal{P}}_{i}$ contains a reasonably large number (set to be 50 in our implementation) of parent sets, we switch to warm-start initialization for pricing at node $i$. This hybrid approach focuses on local refinement around the current best pattern while building upon the foundation of global exploration.
@@ 5 Numerical Experiments
To evaluate the efficiency of our method, we conducted numerical experiments comparing the performance of our Column Generation with DCA-pricing (CG-DCA) method against five baseline methods: GOBNILP, cussens2020gobnilp , CG-MINLP cussens2023branch (in Section 5.1), HC heckerman1995learning (in Appendix D), stable-PC colombo2014order and MMHC tsamardinos2006max . While CG-MINLP is evaluated on Gaussian.test dataset in the R package bnlearn scutari2010learning , other methods are tested on both simulated and real-world datasets. The last three methods are tested using bnlearn. All IP-related experiments are conducted on a Linux machine with two Intel XEON Platinum 8575C processors, with the number of threads used limited to 1. All IPs are solved using Gurobi 11.0.3.
@@ 5.1 Comparing DCA with MINLP for Pricing
For larger instances, CG-MINLP usually fails to produce a valid solution DAG within a reasonable time limit as solving the pricing problem exactly as an MINLP is quite cumbersome in practice. For a graph with node size $n=20$ and average in-degree $d=1$, solving the first pricing problem using MINLP takes an average of $35.1$ seconds per node, while the DCA method takes $0.3$ second per node for the same instance. The difference is even larger for later pricing problems. In addition, MINLP approaches suffer from numerical issues as they require an approximation for the nonlinear $\log$ function (at least in Gurobi), producing inaccurrate estimates for the pricing objective.
@@ 5.2 Results on Simulated Gaussian Data
Graph BIC Score Gap (%) Time (seconds)
$n$ $N$ $d$ CG-DCA GOBNILP CG-DCA GOBNILP
15 5000 0.5 0.00 0.14 0005.80 0008.33
15 5000 1 0.03 0.22 0024.92 0019.74
15 5000 1.5 0.07 0.62 0070.61 0060.09
15 5000 2 0.27 2.06 217.87 167.41
20 5000 0.5 0.00 0.02 0012.43 0003.38
20 5000 1 0.02 0.35 0086.05 287.87
20 5000 1.5 0.22 1.63 0403.38 962.50
20 5000 2 0.65 4.77 1385.88 4603.81
25 5000 0.5 0.00 0.12 0037.28 0021.95
25 5000 1 0.06 0.56 0272.19 1964.11
25 5000 1.5 0.26 1.92 1426.39 5805.01
25 5000 2 0.69 5.15 4850.47 8500.77
@@
Now we test on larger instances with Gaussian data. Following the experimental setup in cussens2023branch ; bhattacharya2021differentiable , we randomly generate Bayesian networks with node size $n\in\{15,20,25\}$, and simulate Gaussian data from those networks. To generate the DAGs, we first fix a topological order of the nodes and then simulate edges according to the specified probabilities where edges must be directed from lower-order nodes to higher-order nodes (there are $n(n-1)/2$ such possible edges). While we do not enforce a maximum in-degree of each node, we randomly simulate edges using their existence probability $p$ determining the overall density of the graph, i.e., for each possible edge we include it in the network with probability $p$. The average in-degree $d$ of nodes in the ground truth DAG is controlled by the edge existence probability since $p={2d}/{(n-1)}$. We vary $p$ such that $d$ is varied among $\{0.5,1,1.5,2\}$.
After generating the random graph structure, we simulate Gaussian data based on the network topology. The linear coefficients for the features were randomly drawn from the interval ±$\pm[0.5,2]$, with additive noise following the Normal distribution $N(0,\sigma^{2})$. The true value of variance parameter $\sigma^{2}$ is uniformly drawn from the interval $[0.7,1.2]$.
Tables 1 and 2 present a comparison between CG-DCA and baseline methods across various node sizes ($n$), and average in-degrees ($d$) with $N=5{,}000$. Results for $N=20{,}000$ exhibit little variation from those for $N=5{,}000$ and can be found in Appendix D. For each graph structure, we generate $10$ independent data instances using different random seeds and record the average performance metrics.
Table 2: Performance Comparison of CG-DCA, GOBNILP (GI), PC and MMHC on Gaussian Datasets
$n,N,d$ Precision Recall SHD
CG-DCA GI PC MMHC CG-DCA GI PC MMHC CG-DCA GI PC MMHC
(15, 5000, 0.5) 0.71 0.66 0.59 0.50 0.79 0.73 0.62 0.51 02.60 02.90 03.00 03.50
(15, 5000, 1) 0.78 0.67 0.72 0.59 0.84 0.72 0.67 0.54 04.10 05.00 04.80 06.70
(15, 5000, 1.5) 0.75 0.73 0.61 0.56 0.82 0.80 0.54 0.48 06.80 07.40 11.50 12.20
(15, 5000, 2) 0.61 0.58 0.54 0.52 0.74 0.65 0.39 0.36 17.90 16.90 20.90 20.40
(20, 5000, 0.5) 0.74 0.71 0.76 0.59 0.77 0.75 0.79 0.60 02.90 03.50 02.60 04.40
(20, 5000, 1) 0.72 0.69 0.75 0.65 0.79 0.77 0.71 0.59 07.00 07.50 05.90 07.90
(20, 5000, 1.5) 0.61 0.61 0.62 0.65 0.79 0.71 0.50 0.49 17.50 16.80 17.90 16.80
(20, 5000, 2) 0.58 0.46 0.54 0.55 0.77 0.53 0.34 0.31 30.00 33.60 31.50 31.10
(25, 5000, 0.5) 0.74 0.61 0.84 0.69 0.79 0.67 0.84 0.69 03.10 05.90 02.20 04.40
(25, 5000, 1) 0.78 0.69 0.80 0.74 0.86 0.76 0.75 0.69 09.20 11.10 07.40 08.70
(25, 5000, 1.5) 0.63 0.57 0.70 0.71 0.82 0.68 0.58 0.56 27.50 22.10 18.50 18.00
(25, 5000, 2) 0.61 0.47 0.53 0.59 0.78 0.54 0.35 0.35 40.30 40.30 38.70 36.50
*Each value is the average over ten independent instances.
In Table 2, we compare four methods with the following metrics on the quality of the solution graphs: (1) precision and recall tillman2011learning , which measure the proportion of true edges among predicted edges and predicted edges among true edges, respectively; (2) the structural Hamming distance (SHD) tsamardinos2006max , which quantifies the dissimilarity between two graphs by counting the required edge additions, deletions, or reversals to make one graph identical to another andersson1997characterization . Prior to computing these metrics, we convert both predicted and true graphs into their corresponding essential graphs andersson1997characterization , which represent their Markov equivalence classes. The conversion of essential graphs is performed following the methodology described in meek2013causal . We also have results for HC but only present them in Appendix D due to its inferior performance. Notably, CG-DCA almost outperforms all methods in all metrics on smaller graphs with $n=15$. CG-DCA achieves better average recall across most instances while maintaining competitive precision and SHD compared to other constraint-based and hybrid approaches. Better performance on recall than on precision and SHD indicates that CG-DCA tends to select denser graphs than others.
5.3 Result on Discrete Data
Table 3: Performance Comparison of CG-DCA, GOBNILP (GI), PC, and MMHC on Discrete Datasets

### v2

@@ 4 Row and Column Generation Scheme
4:  Initialize $\texttt{rc}_{i}\leftarrow$-$\infty$ for $i=1,\ldots,n$ $\triangleright$ initialize reduced costs
@@
10:      $(\texttt{pa},\texttt{rc}_{i})\leftarrow\text{DCA-Pricing}(D,i,x^{0},\lambda^{*},\Lambda,\epsilon)$ $\triangleright$ DCA pricing
11:      $\hat{\mathcal{P}}_{i}\leftarrow\hat{\mathcal{P}}_{i}\cup\{\texttt{pa}\}$ $\triangleright$ column generation
@@
19:     $^{C}\hat{\mathcal{C}}\leftarrow\hat$$\mathcal$$\cup\{C^{*}\}$
@@
22:  $G\leftarrow$ optimal DAG chosen from $\hat{\mathcal{P}}_{1:n}$; simultaneously update $\hat{\mathcal{C}} \triangleright$ row generation for RMIP
@@
Algorithm 2 demonstrates how we incorporate our DCA pricing method into a simultaneous row and column generation framework. The algorithm begins by initializing the candidate parent sets $\hat{\mathcal{P}}_{1:n}$ and the cluster set $\hat{\mathcal{C}}$ with basic candidates (Line 2). It then iterates among three main phases: (1) a CG phase (Lines 4-14) where the RMLP is solved to obtain optimal dual solutions $\lambda^{*}$, followed by the solution of the pricing problem for each node via DCA — The pricing solutions that have negative reduced costs are added to $\hat{\mathcal{P}}_{i}$. The CG phase repeats until all reduced costs become non-negative. In Line 8, the pricing problem on node $i$ is solved only when the current reduced cost $rc_{i}$ is negative, to avoid spending too much time on pricing; (2) a row generation phase (Lines 15-21) where the RMLP optimal solution is used to identify the most violated cluster constraint (see Appendix B for details) which is added to cluster set $\hat{\mathcal{C}}$; (3) the integer solution phase (Line 22), where the algorithm finds the optimal DAG $G$ with current candidate $\hat{\mathcal{P}}_{1:n}$ — This is achieved by solving RMIP using branch and cut with callbacks (a feature that allows users to interrupt the solution process in modern IP solvers such as Gurobi), where we check if any cluster constraint is violated when we meet a integer incumbent and add violated cluster constraints as lazy constraints. The three phases iterate until convergence to a valid DAG solution, progressively refining both the columns in $\hat{\mathcal{P}}_{1:n}$ and cluster constraints in $\hat{\mathcal{C}}$.
When we implement the CG phase in Algorithm 2, in addition to the final solution generated by DCA, we add to $\hat{\mathcal{P}}_{i}$ all candidate parent sets associated with intermediate $x^{t}$ solutions that have negative reduced costs. Empirically, we find that it helps accelerate convergence and improves solution quality. In addition, we use the warmstart initialization in Line 22 for the master IP problem to quickly drive up the primal bound.
@@ 4.1 Choice of the Initial Point
A step that remains ambiguous in Algorithm 2 is the choice of the initial point $x^{0}$ for DCA (Line 9), which we find having a significant impact on the overall solution time and solution quality. Below we first describe two initialization strategies.
The first strategy is a warm-start initialization strategy. Note that by the complementarity property of optimal solutions in LP [8], the parent sets selected (i.e., parent sets $J\in\hat{\mathcal{P}}_{i}$ associated with the primal variables $x_{i\leftarrow J}$ taking positive values) in the RMLP solution have reduced costs equal to zero. Therefore, we consider using the parent sets selected in the RMLP solution to construct the initial solution $x^{0}$ for DCA. Due to the monotonicity property (4) of DCA, it often quickly finds columns with negative reduced costs, leading to a rapid convergence of DCA. However, a drawback of the warm-start initialization strategy is that it may restrict exploration of parent sets as the evaluation remains localized around columns already contained in the candidate parent set choice set. Preliminary experiments show that a pure warm-start initialization strategy can quickly find a DAG with a good score while the solution often deviates a lot from the ground truth due to its greedy nature.
The second strategy employs a random initialization that samples a point $x^{0}$ uniformly from $0,1]^{V\setminus\{i\}}$. This random initialization strategy facilitates global exploration on the parent set space and may discover patterns distant from the current ones in $\hat{\mathcal{P}}_{i}$. However, we find that exclusive reliance on random initialization can significantly slow down DCA convergence, as computational resources may be expended on evaluating solutions far from the optimum.
A hybrid initialization strategy that combines efficiency and diversity of the previous two strategies is often more effective in practice. In our implementation, we initially employ random initialization to broadly explore the solution space. Once the candidate set $\hat{\mathcal{P}}_{i}$ contains a reasonably large number (set to be 50 in our implementation) of parent sets, we switch to warm-start initialization for pricing at node $i$. This hybrid approach focuses on local refinement around the current best pattern while building upon the foundation of global exploration. A comparison of the three initialization strategies can be found in Appendix D.1.
@@ 5 Numerical Experiments
To evaluate the efficiency of our method, we conducted numerical experiments comparing the performance of our Column Generation with DCA-pricing (CG-DCA) method against five baseline methods: GOBNILP [17], CG-MINLP [18] (in Section 5.1), HC [28] (in Appendix D.2), stable-PC [14] and MMHC [47]. For score-based approaches, we use the BIC score as our scoring function. While CG-MINLP is evaluated on Gaussian.test dataset in the R package bnlearn [43], other methods are tested on both simulated and real-world datasets. We implement CG-DCA in python, and utilize the python implementation of GOBNILP. The last three baseline methods are tested using bnlearn in R. All IP-related experiments are conducted on a Linux machine with two Intel XEON Platinum 8575C processors, with the number of threads used limited to 1. All IPs are solved using Gurobi 11.0.3.
@@ 5.1 Comparing DCA with MINLP for Pricing
For larger instances, CG-MINLP usually fails to produce a valid solution DAG within a reasonable time limit as solving the pricing problem exactly as an MINLP is quite cumbersome in practice. For a graph with node size $n=20$ and average in-degree $d=1$ (N=5, [^0] N=5{,}000, simulated as in Section 5.2), solving the first pricing problem using MINLP takes an average of $35.1$ seconds per node, while the DCA method takes $0.3$ second per node for the same instance. The difference is even larger for later pricing problems. In addition, MINLP approaches suffer from numerical issues as they require an approximation for the nonlinear $\log$ function (at least in Gurobi), producing inaccurrate estimates for the pricing objective.
@@ 5.2 Results on Simulated Gaussian Data
$(n,N,d)$ BIC Score Gap (%) Time (seconds)
CG-DCA GOBNILP CG-DCA GOBNILP
(15, 05000, 0.5) (0.00, 0.00) [0.00] (0.00, 01.46) [0.15] (001.12, 00014.71) [0004.06] (0005.47, 00024.28) [0008.33]
(15, 05000, 1.0) (0.00, 0.24) [0.03] (0.00, 01.06) [0.22] (002.54, 00093.58) [0025.76] (0005.72, 00056.11) [0019.74]
(15, 05000, 1.5) (0.00, 0.62) [0.16] (0.00, 02.18) [0.62] (017.95, 00196.01) [0076.72] (0011.51, 00137.52) [0060.09]
(15, 05000, 2.0) (0.00, 0.84) [0.39] (0.48, 04.88) [2.06] (041.13, 00340.22) [0170.39] (0022.99, 00641.90) [0167.41]
(15, 20000, 0.5) (0.00, 0.00) [0.00] (0.00, 00.64) [0.06] (003.53, 00035.11) [0010.17] (0005.53, 00012.62) [0007.35]
(15, 20000, 1.0) (0.00, 0.27) [0.03] (0.00, 01.19) [0.23] (007.52, 00100.20) [0036.71] (0001.31, 00093.02) [0019.67]
(15, 20000, 1.5) (0.00, 0.11) [0.02] (0.00, 03.10) [0.86] (025.17, 00265.33) [0104.25] (0012.34, 00285.39) [0076.29]
(15, 20000, 2.0) (0.00, 0.94) [0.23] (0.98, 03.97) [2.01] (055.11, 00335.96) [0229.80] (0011.66, 00190.19) [0089.92]
(20, 05000, 0.5) (0.00, 0.01) [0.00] (0.00, 00.18) [0.02] (002.49, 00020.20) [0009.53] (0001.36, 00006.99) [0003.38]
(20, 05000, 1.0) (0.00, 0.08) [0.01] (0.00, 01.38) [0.35] (005.97, 00150.18) [0084.34] (0001.59, 01714.57) [0287.87]
(20, 05000, 1.5) (0.01, 0.51) [0.22] (0.00, 04.30) [1.63] (144.94, 00539.72) [0312.67] (0121.01, 03483.96) [0962.50]
(20, 05000, 2.0) (0.22, 1.15) [0.70] (0.07, 10.42) [4.77] (403.98, 01904.24) [1039.53] (1063.89, 08703.57) [4603.81]
(20, 20000, 0.5) (0.00, 0.00) [0.00] (0.00, 00.35) [0.04] (006.57, 00036.37) [0018.42] (0001.50, 00011.25) [0004.92]
(20, 20000, 1.0) (0.00, 0.22) [0.03] (0.00, 01.49) [0.22] (013.72, 00330.05) [0133.69] (0001.54, 01371.66) [0261.89]
(20, 20000, 1.5) (0.00, 0.43) [0.15] (0.00, 03.17) [1.45] (216.26, 00821.01) [0425.17] (0135.59, 07269.60)[2457.26]
(20, 20000, 2.0) (0.09, 1.08) [0.54] (0.21, 07.05) [3.63] (457.33, 01739.24) [0997.99] (0517.67, 10801.59) [4394.45]
(25, 05000, 0.5) (0.00, 0.00) [0.00] (0.00, 00.82) [0.12] (002.73, 00067.05) [0021.97] (0001.61, 00151.24) [0021.95]
(25, 05000, 1.0) (0.00, 0.25) [0.03] (0.00, 01.83) [0.56] (028.04, 00953.40) [0261.84] (0011.14, 07451.37) [1964.11]
(25, 05000, 1.5) (0.00, 1.11) [0.24] (0.09, 05.27) [1.92] (144.19, 02880.93) [0949.90] (0084.46, 10801.61) [5805.01]
(25, 05000, 2.0) (0.00, 2.18) [0.86] (1.84, 10.18) [5.15] (448.93, 12665.38) [3054.63] (0723.38, 10801.63) [8500.77]
(25, 20000, 0.5) (0.00, 0.01) [0.00] (0.00, 00.78) [0.09] (009.50, 00164.57) [0050.22] (0001.62, 00169.94) [0023.96]
(25, 20000, 1.0) (0.00, 0.22) [0.06] (0.00, 02.72) [0.40] (070.34, 01287.17) [0404.43] (0011.86, 10801.48) [1948.05]
(25, 20000, 1.5) (0.00, 1.19) [0.23] (0.66, 05.47) [2.21] (237.86, 04955.58) [1372.59] (0115.63, 10801.73) [5853.91]
(25, 20000, 2.0) (0.00, 2.71) [0.67] (1.50, 08.30) [4.51] (733.04, 07371.52) [3189.98] (2302.81, 10801.71) [7967.28]
*The values are the (min, max)[average] over ten independent instances.
Now we test on larger instances with Gaussian data. Following the experimental setup in [18, 9], we randomly generate Bayesian networks with node size $n\in\{15,20,25\}$, and simulate Gaussian data from those networks. To generate the DAGs, we first fix a topological order of the nodes and then simulate edges according to the specified probabilities where edges must be directed from lower-order nodes to higher-order nodes (there are n(n1)/2n(n-1)/2 such possible edges). While we do not enforce a maximum in-degree of each node, we randomly simulate edges using their existence probability $p$ determining the overall density of the graph, i.e., for each possible edge we include it in the network with probability $p$. The average in-degree $d$ of nodes in the ground truth DAG is controlled by the edge existence probability since p=2d/(n1)p={2d}/{(n-1)}. We vary $p$ such that $d$ is varied among $\{0.5,1,1.5,2\}$.
After generating the random graph structure, we simulate Gaussian data based on the network topology. The linear coefficients for the features were randomly drawn from the interval $\pm[0.5,2]$, with additive noise following the Normal distribution $N(0,\sigma^{2})$. The true value of variance parameter $\sigma^{2}$ is uniformly drawn from the interval $[0.7,1.2]$.
Tables 1 and 2 present a comparison between CG-DCA and baseline methods across various node sizes ($n$), and average in-degrees ($d$) with N=5, [^0] N=5{,}000 and N=20, [^0] N=20{,}000. For each graph structure, we generate $10$ independent data instances using different random seeds and record the average performance metrics.
Table 2: Performance Comparison of CG-DCA, GOBNILP (GI), (Stable-)PC and MMHC on Gaussian Datasets
$(n,N,d)$ Precision Recall SHD
CG-DCA GI PC MMHC CG-DCA GI PC MMHC CG-DCA GI PC MMHC
(15, 05000, 0.5) 0.78 0.66 0.59 0.50 0.80 0.73 0.62 0.51 01.70 02.90 03.00 03.50
(15, 05000, 1.0) 0.70 0.67 0.72 0.59 0.78 0.72 0.67 0.54 06.20 05.00 04.80 06.70
(15, 05000, 1.5) 0.65 0.73 0.61 0.56 0.77 0.80 0.54 0.48 11.80 07.40 11.50 12.20
(15, 05000, 2.0) 0.55 0.58 0.54 0.52 0.72 0.65 0.39 0.36 21.90 16.90 20.90 20.40
(15, 20000, 0.5) 0.71 0.69 0.68 0.53 0.74 0.73 0.68 0.53 02.20 02.20 01.90 03.00
(15, 20000, 1.0) 0.76 0.78 0.71 0.61 0.78 0.81 0.69 0.59 04.10 04.00 04.70 06.20
(15, 20000, 1.5) 0.76 0.73 0.72 0.66 0.85 0.79 0.66 0.57 06.90 09.40 08.60 10.00
(15, 20000, 2.0) 0.61 0.53 0.53 0.55 0.75 0.60 0.43 0.39 20.60 19.90 20.01 19.80
(20, 05000, 0.5) 0.74 0.71 0.76 0.59 0.80 0.75 0.79 0.60 03.40 03.50 02.60 04.40
(20, 05000, 1.0) 0.71 0.69 0.75 0.65 0.76 0.77 0.71 0.59 03.70 07.50 05.90 07.90
(20, 05000, 1.5) 0.50 0.61 0.62 0.65 0.71 0.71 0.50 0.49 28.20 16.80 17.90 16.80
(20, 05000, 2.0) 0.44 0.46 0.54 0.55 0.66 0.53 0.34 0.31 36.40 33.60 31.50 31.10
(20, 20000, 0.5) 0.82 0.80 0.73 0.58 0.83 0.82 0.76 0.59 02.20 02.60 03.00 04.60
(20, 20000, 1.0) 0.61 0.73 0.79 0.68 0.70 0.79 0.77 0.65 08.30 06.90 04.90 07.10
(20, 20000, 1.5) 0.65 0.58 0.64 0.63 0.81 0.68 0.56 0.50 16.70 18.20 16.00 16.50
(20, 20000, 2.0) 0.52 0.48 0.51 0.56 0.73 0.55 0.37 0.33 33.00 32.40 31.30 29.30
(25, 05000, 0.5) 0.82 0.61 0.84 0.69 0.84 0.67 0.84 0.69 02.70 05.90 02.70 04.40
(25, 05000, 1.0) 0.81 0.69 0.80 0.74 0.86 0.76 0.75 0.69 06.70 11.10 07.40 08.70
(25, 05000, 1.5) 0.66 0.57 0.70 0.71 0.75 0.68 0.58 0.56 13.60 22.10 18.50 18.00
(25, 05000, 2.0) 0.54 0.47 0.53 0.59 0.73 0.54 0.35 0.35 45.70 40.30 38.70 36.50
(25, 20000, 0.5) 0.76 0.80 0.81 0.70 0.80 0.84 0.83 0.70 03.70 02.80 02.80 04.30
(25, 20000, 1.0) 0.76 0.71 0.81 0.68 0.86 0.76 0.78 0.64 09.10 09.90 06.70 09.90
(25, 20000, 1.5) 0.57 0.53 0.73 0.63 0.67 0.66 0.63 0.51 14.10 25.00 16.50 19.50
(25, 20000, 2.0) 0.54 0.48 0.61 0.59 0.66 0.56 0.46 0.40 23.90 39.10 33.80 34.10
@@
In Table 2, we compare four methods with the following metrics on the quality of the solution graphs: (1) precision and recall [46], which measure the proportion of true edges among predicted edges and predicted edges among true edges, respectively; (2) the structural Hamming distance (SHD) [47], which quantifies the dissimilarity between two graphs by counting the required edge additions, deletions, or reversals to make one graph identical to another [3]. Prior to computing these metrics, we convert both predicted and true graphs into their corresponding essential graphs [3], which represent their Markov equivalence classes. The conversion of essential graphs is performed following the methodology described in [38]. We also have results for HC but only present them in Appendix D.2 due to its inferior performance. Notably, CG-DCA achieves better average recall across most instances while maintaining competitive precision and SHD compared to other constraint-based and hybrid approaches. Better performance on recall than on precision and SHD indicates that CG-DCA tends to select denser graphs than others.
5.3 Results on Discrete Data
Table 3: Performance Comparison of CG-DCA, GOBNILP (GI), (Stable-)PC, and MMHC on Discrete Datasets